OpenStudy (anonymous):

How would I go about finding which quadrant theta/2 is in, given that theta is a quadrant 3 angle?

OpenStudy (zzr0ck3r):

theta in quadrent two implies \(\pi \le \theta \le \frac{3\pi}{2}\implies\frac{\pi}{2}\le\frac{\theta}{2}\le\frac{3\pi}{4}\)